How many years would it take you to count and Avogadro’s number of objects if you could count at the rate of five objects per second?
1
Expert's answer
2020-10-28T08:43:46-0400

Avogadro's number is "6.022 \times 10^{23}"


So the number of objects you want to count is "6.022 \times 10^{23}"


If you count them at a rate of 5 objects per second, it'd take;

"\dfrac{6.02 \times 10^{23} \textsf{ objects}}{5 \textsf{ objects per second}}"


"=1.2046 × 10^{23}\, s"



"1 \textsf{ year} = 60 \textsf{ seconds per minute} × 60 \textsf{ minutes per hour} × 24 \textsf{ hours per day} × 365 \textsf{ days per year}\\
=31536000 \textsf{ seconds}\\
=3.1536 × 10^7 \textsf{ seconds}"


"\therefore \textsf{number of years} = \dfrac{1.2046 × 10^{23}\, s}{3.1536 × 10^7 \textsf{ seconds per year}}"

"= 3.82 × 10^{15} \textsf{ years}"


"= 382,000,000,000,000 \textsf{ years}"


"= 382,000,000,000,000 \textsf{ trillion years}"